A boy is 3 times as old as a girl. If their total age is 48years, how old is the girl

Answer by macston(5194)

.
G=girl's age; B=boy's age=3G
.
G+B=48 years . Substitute for B.
G+3G=48 years
4G=48 years
G=12 years
.
ANSWER : The girl is 12 years old.
.
CHECK:
B=3G=3(12)=36 years
The boy is 36 years old.
.
G+B=48 years
12 years + 36 years = 48 years
48 years = 48 years
.